Three women go camping together and are setting up camp. First, they decide to set up a large tarp and they open it up. One woman grabs the Northeast corner and pulls with a force of 18 N 60° North of East. Another woman grabs the Southeast corner and pulls it back at an angle of 40° South of East. The third woman grabs in the middle of the West side of the tarp and pulls straight back in the West direction. What are the magnitudes of the pulling forces that the last two women mentioned apply to the tarp, if it unfolds nicely but the centre but remains stationary?


1
Expert's answer
2021-01-10T18:29:53-0500
"F_2 \sin{40}=18 \sin{60}\\F_2=24.25\ N"

"F_3=24.24 \cos{40}+18 \cos{60}\\F_3=27.57\ N"